A given volume of a gas G diffuses through amemebrane in 10s. Under the same condition of temperature and pressure, an equal volume of oxygen diffuses through he same membrane in 12.5 seconds. Determine the molecular mass of gas G
1
Expert's answer
2020-05-05T03:37:32-0400

R1/R2=√M2/M1.

M2=(R1/R2)^2 x M1= (12.5/10)^2 x 32 = 50.